OpenStudy (anonymous):

the reciprocal 4/11 is?

OpenStudy (anonymous):

11/4

OpenStudy (anonymous):

what he said.

OpenStudy (anonymous):

11/4

OpenStudy (anonymous):

11/4

OpenStudy (anonymous):

1 / 4/11 = 11/4

OpenStudy (anonymous):

11/4 = 2 3/4

OpenStudy (anonymous):

appericate it

OpenStudy (anonymous):

no probs